OpenStudy (anonymous):

Helpppp What Polynomial has roots of -6,1 and 4? x^3-9x^2-22x+24 x^3-x^2-26x-24 x^3+x^2-26x+24 x^3+9x+14x-24

OpenStudy (anonymous):

(x-6)(x+1)(x+4)=x^3-x^2-26x-24

OpenStudy (anonymous):

In order for those numbers, -6, 1 and 4 to be roots the function must equal to zero when x=-6, x=1 and x=4. Plugging in is faster than factoring, the answer is x^3+x^2-26x+24

OpenStudy (anonymous):

which one is correct??

OpenStudy (anonymous):

damoss is correct, i should have written (x+6)(x-1)(x-4)=x^3+x^2-26 x+24

OpenStudy (anonymous):

you need to flip the sign of the numbers first before factoring them

OpenStudy (anonymous):

did you get my message

OpenStudy (anonymous):

What is the simplified form of 4x plus 2 over x plus 5 minus the fraction 3x plus 1 over x plus 5 ?

OpenStudy (anonymous):

\[\large \frac{ 4x+2 }{ x+5 }-\frac{ 3x+1 }{x+5? }\]

OpenStudy (anonymous):

\[\large \frac{ 4x+2-(3x+1) }{ x+5 }\]

OpenStudy (anonymous):

\[\frac{ x+1 }{ x+5 }\]